52.7 ÷ _______ = 0.527 
 52.7 ÷ _______ = 0.527 
 52.7 ÷ _______ = 0.527 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

52.7 ÷ 100 = 0.527
Let us assume the missing fraction be x, Then,

52.7 ÷ x = 0.527

52.7/x = 0.527
By cross multiplication,
we get
x = 52.7/0.527
x = 52700/527 ... [divide both numerator and denominator by 527]
x = 100




Thanks